let's get some real life answers. in dog racing there are 8 dogs in each race. to cash a quinela bet you need to pick 2 dogs and have them finish first and second (order not important). if you were to pick 3 of the 8 dogs, what is the probability that two of them will finish first and second?
1
Expert's answer
2012-07-27T11:11:50-0400
P = 3 * 1/8 + 2 * 1/7 = 6/56 = approx. 10.71%